OpenStudy (hihi67):

@A water drop has a volume of 0.08 cm 3 Give all answers to using powers of ten! -How many drops of water give 10 liters? -How many drops fill a bathtub (200 liters)? Pro-days to evaporate around the world 1 1170 km 3 of water. How many liters are they? How many drops are these? @A water drop has a volume of 0.08 cm 3 Give all answers to using powers of ten! -How many drops of water give 10 liters? -How many drops fill a bathtub (200 liters)? Pro-days to evaporate around the world 1 1170 km 3 of water. How many liters are they? How many drops are these? @Mathematics

OpenStudy (hihi67):

its cm ^3

OpenStudy (hihi67):

please please help!

OpenStudy (mertsj):

Since 1 cm^3 = 1 ml and 1000 ml = 1 liter\[10 liters(1000ml/1liter)(1cm^3/1ml)(1drop water/.08cm^3)=125000=1.25X10^5\]

OpenStudy (hihi67):

howd you get 125000?

OpenStudy (mertsj):

Write my solution on paper using straight fraction bars. Cancel the units just as if they were numbers and then multiply the numbers on top and divide by the numbers on the bottom.

OpenStudy (mertsj):

You end up with 10(1000) divided by .08

OpenStudy (hihi67):

does it matter that the cm is cubed?

OpenStudy (mertsj):

cm^3 cancel cm^3 so you end up with drops of water

OpenStudy (hihi67):

ok... and for the second question. is it just 20000ml divided by .08?

OpenStudy (mertsj):

200 liters = 200(1000) ml so it is 200,000 ml

OpenStudy (hihi67):

right.

OpenStudy (hihi67):

and the 3rd one?

OpenStudy (mertsj):

I haven't a clue what that means. What are pro days?

OpenStudy (hihi67):

1 1170 km=11170000ml?

OpenStudy (hihi67):

divided by 0.08

OpenStudy (anonymous):

no much much more.... 1Km^3=10^(12) Liters

OpenStudy (mertsj):

What are pro-days?

OpenStudy (hihi67):

how so?

OpenStudy (mertsj):

Please read what you posted and explain what it means. "Pro-days to evaporate around the world 1 1170 km 3 of water"

OpenStudy (hihi67):

no idea.

OpenStudy (mertsj):

Then where did you get it?

OpenStudy (hihi67):

i tranlated it.

OpenStudy (mertsj):

Better translate it again. Something is wrong.

OpenStudy (hihi67):

i think it means in the whole world 1 1170 ^3 water evaporates

OpenStudy (anonymous):

1L=1dm^3 and 1Km^3=10^(12)dm^3 so 1Km^3=10^12 L=10^15 ml and 11170 Km^3=11170x10^(15) ml this gives 139625x10^(15) drops...

OpenStudy (hihi67):

how is 1L 1dm^3?

OpenStudy (anonymous):

it is so by definition,they are both units of volume

OpenStudy (hihi67):

oh

OpenStudy (hihi67):

how is 1Km^3=10^(12) Liters?

OpenStudy (hihi67):

does it have something to do with the cubed?

OpenStudy (anonymous):

yes 1km=10^4 dm so 1km^3=(10^4)^3 dm^3=10^12 dm^3=10^12 L

OpenStudy (hihi67):

how did you get liters? km is not volume.

OpenStudy (anonymous):

no but km^3 is

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Litre

OpenStudy (hihi67):

oooooooh

OpenStudy (anonymous):

@ Mertsj 1KM^3=100000^3 CM^3

OpenStudy (hihi67):

how does 1 cm^3 = 1 ml ?

OpenStudy (mertsj):

1km^3 = 1000^3 m^3 = 1000^3(100^3)cm^3 = 10^9(10^6)cm^3 = 10^15 cm^3=10^15 ml = 10^12 liters

OpenStudy (hihi67):

huh? how does 1 cm^3 = 1 ml ?

OpenStudy (mertsj):

Well to demonstrate that you would have to get a little container that is a 1 cm by 1cm by 1 cm, fill it with water and pour it into a graduated cylinder. It would come up to the 1 ml mark and then if you weighed the water, it would weigh 1 gram.

OpenStudy (hihi67):

what about 1 cm and 1 ml?

OpenStudy (hihi67):

can you not do that?

OpenStudy (anonymous):

u can't compare a unit of length with a unit of volume...keep this in mind

OpenStudy (mertsj):

But you can compare a cubed unit of length with a unit of volume

OpenStudy (mertsj):

Because now it is volume